find the ordered pair
y=3 x -9
1) (3,_)
2) (4,_)
3) (0,_)

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

find the ordered pair
y=3x-9
---------
1) (3,_)
y = 3*3-9 = 0
---------------------
2) (4,_)
y = 3*4-9 = 3
---------------------
3) (0,_)
y = 3*0-3 = -3
========================
Cheers,
Stan H.

Answer by rfer(16322)

1) (3,0)
2) (4,4)
3) (0,-9)